# Efficient Tactile Perception with Soft Electrical Impedance Tomography and Pre-trained Transformer

### v1

@@ Abstract
Tactile sensing is fundamental to robotic systems, enabling interactions through physical contact in multiple tasks. Despite its importance, achieving high-resolution, large-area tactile sensing remains challenging. Electrical Impedance Tomography (EIT) has emerged as a promising approach for large-area, distributed tactile sensing with minimal electrode requirements which can lend itself to addressing complex contact problems in robotics. However, existing EIT-based tactile reconstruction methods often suffer from high computational costs or depend on extensive annotated simulation datasets, hindering its viability in real-world settings. To address this shortcoming, here we propose a Pre-trained Transformer for EIT-based Tactile Reconstruction (PTET), a learning-based framework that bridges the simulation-to-reality gap by leveraging self-supervised pretraining on simulation data and fine-tuning with limited real-world data. In simulations, PTET requires 99.44% fewer annotated samples than equivalent state-of-the-art approaches (2,500 vs. 450,000 samples) while achieving reconstruction performance improvements of up to 43.57% under identical data conditions. Fine-tuning with real-world data further enables PTET to overcome discrepancies between simulated and experimental datasets, achieving superior reconstruction and detail recovery in practical scenarios. PTET’s improved reconstruction accuracy, data efficiency, and robustness in real-world tasks establish it as a scalable and practical solution for tactile sensing systems in robotics, especially for object handling and adaptive grasping under varying pressure conditions.
Index Terms:
Masked Autoencoder, Electrical Impedance Tomography, Tactile Sensing, Human-Machine Interaction.
@@ I Introduction
Tactile sensing is a foundational technology in robotics, enabling robots to effectively interact with their environment by perceiving pressure, texture, and force through physical contact [1]. This capability is essential for tasks such as object manipulation, human-robot interaction, and environment exploration [2, 3, 4]. Despite its significance, achieving high-resolution, large-area tactile sensing remains a significant challenge due to inherent trade-offs among accuracy, scalability, and complexity in sensor design. Current tactile sensors typically rely on capacitive [5], piezoresistive [6], or piezoelectric [7, 8] mechanisms. However, the spatial resolution of these systems is constrained by the number of sensor array pixels, which require densely packed sensor arrays for higher spatial resolution. This not only complicates fabrication and increases costs but also limits flexibility and scalability for expanded tactile sensing tasks [9].
@@
To overcome these limitations, Electrical Impedance Tomography (EIT) has emerged as a promising alternative [10]. EIT offers several advantages over traditional methods, including the ability to achieve large-area tactile sensing with fewer boundary electrodes [11, 12, 13]. By injecting small electrical currents through electrode pairs and measuring the resulting voltage changes, EIT reconstructs the conductivity distribution of a conductive soft layer to infer tactile inputs such as pressure or deformation [14, 15]. However, EIT tactile reconstruction poses significant challenges in terms of accuracy and computational efficiency [16].
Model-based methods [17, 18, 19, 20] have been typically employed for EIT reconstruction, solving the inverse problem non-iteratively or iteratively to estimate the conductivity distribution from voltage measurements. While these approaches do not require pre-collected datasets, they often produce suboptimal reconstruction results, especially for dynamic tactile sensing tasks. This limitation is largely attributed to EIT’s inherently ill-posed and ill-conditioned inverse problem [16], exacerbating reconstruction quality issues. Furthermore, the accuracy of model-based approaches is inherently limited due to simplified assumptions on the governing physical model, sensor geometry and materials properties [21].
In recent years, data-driven approaches, particularly those based on deep learning, have gained traction in the EIT domain [22, 23, 24]. These methods aim to train deep nets to directly map voltage measurements to tactile reconstructions, bypassing the iterative processes and linear approximations. While data-driven approaches have shown promise in delivering faster and more accurate reconstructions, they are heavily dependent on large, annotated datasets [25, 26, 27, 28]. Acquiring extensive real-world datasets with ground-truth tactile annotations ($e.g.,>100k$) is often impractical due to the labour-intensive and time-consuming nature of data collection. Consequently, these methods frequently depend on simulation-generated data for model training. However, the inherent discrepancies between simulation and experimental data can introduce considerable reconstruction errors in practical applications. These discrepancies manifest themselves as reconstruction errors, including inaccuracies in pressure magnitude, the appearance of spurious residual artefacts, and misrepresentation of the shape of the contact interface in complex scenarios [29, 30].
To address this issue, self-supervised pretraining has emerged as a powerful technique, enabling models to extract general features from unlabelled data and subsequently transfer the features to various downstream tasks by supervised fine-tuning using limited labelled data. This paradigm has demonstrated success in robotics applications, such as control [31], navigation [32], visual perception [33], and sim-to-real transfer [34]. Recent advancements in self-supervised learning have particularly focused on masked prediction in large models like GPT [35], and BERT [36], showcasing their potential in stacking large language/vision models. For high-dimensional data, one pivotal work in computer vision is the masked autoencoder (MAE) [37], which conducts a scalable vision learner to extract image features. Inspired by this, EIT can similarly be pre-trained in an unlabelled manner using a self-supervised approach. This enables EIT tactile reconstruction models to learn from unlabelled data and subsequently fine-tuned with limited paired (labelled) real-world data, establishing a real-to-real learning pipeline that eliminates the need for extensive real-world datasets while mitigating simulation-to-reality discrepancies.
@@
Figure 1: Electrical Impedance Map (EIM). (a) EIT-based tactile sensing principle. (b) Original EIT measurements. (c) Construction of EIM: an example. (d) Enhanced EIM (E2IM).
@@
In this paper, we propose the Pre-trained Transformer for the EIT-based Tactile Reconstruction (PTET) model, a novel framework that addresses two key challenges in learning-based EIT tactile sensing: (1) the heavy dependency on large annotated datasets and (2) the simulation-to-reality gaps that hinder practical applications. PTET leverages a self-supervised pretraining strategy to learn generalizable features from unlabelled tactile sensing data, significantly reducing reliance on large annotated datasets. The model is subsequently fine-tuned using a limited amount of annotated real-world data, enabling more accurate tactile reconstruction while mitigating inaccuracies typically introduced by simulation-based training. This dual-phase learning strategy allows PTET to bridge the gap between experimental conditions and practical applications. Unlike traditional learning-based approaches that require extensive labelled datasets, which are typically generated from simulation, PTET effectively operates with a small fraction of the real-world data, making it scalable and practical for real-world deployments. The main contributions of this work are as follows:
@@
  • We develop a flexible, large-area EIT-based tactile sensor that introduces a dual-conductivity layer structure, uniquely combining hydrogel for flexibility and biocompatibility with a carbon black-to-graphite (CBG) composite for significantly enhanced conductivity and mechanical flexibility. Compared to existing designs, our sensor achieves superior signal stability and high sensitivity, capable of detecting compressive strain as small as 1 mm over a large sensing area.
  • For the first time, we demonstrate high-resolution tactile reconstruction using a limited experimental dataset in a learning-based framework. This real-to-real approach eliminates errors and mismatches associated with simulated data, making EIT-based tactile sensing more practical and robust for real-world applications.
@@ II Enhanced Electrical Impedance Map
In each EIT measurement, two neighbouring electrodes are used for current injection, while the remaining electrodes are selected for voltage measurements. Among the measuring electrodes, measurements are taken between neighbouring electrode pairs, excluding the current injection electrodes (see Fig. 1a). For the EIT-based tactile sensor with 16 electrodes, this results in 16 (excitation electrode pairs) × 13 (measurement electrode pairs) = 208 measurements per frame, though only 104 independent voltages are recorded due to the reciprocity theorem.
To capture the spatial and geometric features of the EIT measurements in an image-like format [38], we convert the original voltage measurements (104 $\times$1 vector, Fig. 1b) to a 16 $\times$ 16 matrix, termed the Electrical Impedance Map (EIM) (Fig. 1c). The EIM matrix is constructed in Algorithm 1.
Algorithm 1 Construction of EIM
1:$\mathbf{v}_{\text{ori}}$: Original voltage measurements (104 $\times$ 1 vector)
2:$\mathbf{v}_{\text{eim}}$: EIM ($16\times 16$ matrix)
3:Initialize $\mathbf{v}_{\text{eim}}\leftarrow\mathbf{0}_{16\times 16}$triangleright Zero matrix of size $16\times 16$
4:$index\leftarrow 1$
5:for $i=1$ to $16$ do
6:     for $j=1$ to $16$ do
7:         if $j>i+1$ and $index\leq 104$ then
8:              $\mathbf{v}_{\text{eim}}(i,j)\leftarrow\mathbf{v}_{\text{ori}}[index]$
9:              $\mathbf{v}_{\text{eim}}(j,i)\leftarrow\mathbf{v}_{\text{ori}}[index]$triangleright Apply reciprocity theorem
10:              $index\leftarrow index+1$
11:         end if
12:     end for
13:end for
14:return $\mathbf{v}_{\text{eim}}$
To further enhance the representation quality of EIM, we propose using Bicubic interpolation to enhance the EIM to the Enhanced-EIM (E2IM) (following the Algorithm 2). Unlike the original $16\times 16$ matrix where each voltage measurement occupies a single pixel, the E2IM represents each measurement as a $4\times 4$ patch, generating a $64\times 64$ matrix (Fig. 1d). This augmentation improves the spatial granularity and feature richness of the original EIM, enhancing the model’s capability to capture fine-grained spatial details and structural patterns, essential for achieving high-resolution tactile reconstructions and improving the robustness of the learning process.
Algorithm 2 Construction of E2IM
1:$\mathbf{v}_{\text{eim}}$: EIM ($16\times 16$ matrix)
2:$\mathbf{v}$: E2IM ($64\times 64$ matrix)
3:Initialize $\mathbf{v}\leftarrow\mathbf{0}_{64\times 64}$triangleright Zero matrix of size $64\times 64$
4:Apply bicubic interpolation to $\mathbf{v}_{\text{eim}}$ to resize it from $16\times 16$ to $64\times 64$
5:Assign the resized matrix to $\mathbf{v}$
6:return $\mathbf{v}$
@@ III PTET: Pre-trained Transformer for EIT-based Tactile Reconstruction
Figure 2: Overview of PTET. Our method comprises three main processes: bottom— 1) Pretraining with E2IM, for an effective and robust voltage encoder $\mathcal{E}_{vol}$, through self-supervised learning by predicting masked regions in grey. 2) Pretraining with tactile map, through self-supervised learning, for a robust tactile decoder $\mathcal{D}_{tact}$; top— 3) Fine-tuning with limited pairs from real-world measurements, the fine-tuned model will be deployed for reconstruction. Notably, during fine-tuning, $\mathcal{E}_{vol}$’s parameters are frozen, while $\mathcal{D}_{tact}$’s are tunable.
The reconstruction of tactile representations t from EIT voltage measurements v, represented in the format of E2IM, can be described by a direct mapping $f$:
$f:\textbf{v}\rightarrow\textbf{t}.$ (1)
@@
    Pretraining with E2IM: In this stage, a masked autoencoder $\mathcal{D}_{vol}(\mathcal{E}_{vol}(\textbf{v}))$ is employed to extract features from unlabeled voltage data v, which is represented by E2IM. This process trains an effective and robust encoder $\mathcal{E}_{vol}(\textbf{v})$, functioning mapping $p_{v}:\textbf{v}\rightarrow\textbf{z}_{v}$, which compresses the input v to a latent variable $\textbf{z}_{v}$;
@@
The PTET framework adopts a dual-phase training scheme: the pretraining phase with 1) and 2) and fine-tuning phase with 3). In the pretraining phase, the functioning $\mathcal{E}_{vol}$ and $\mathcal{D}_{tact}$ are trained with large-scale unannotated datasets. It allows the model to learn generalizable features from E2IM and tactile maps separately based on a large scale of random simulation data, which can be readily acquired. While in the fine-tuning phase, using only limited paired data from real-world measurements, the pretrained models are aligned to establish the relationship between the voltage and tactile domains. The tactile map can then be reconstructed from a given E2IM describing real-world voltage measurements. Notably, in both 1) and 2), PTET employs an asymmetric architecture for encoders and decoders to enhance the efficiency and effectiveness of feature representation in $\mathcal{E}_{vol}$ and $\mathcal{D}_{tact}$, as detailed in the following sections.
@@ III-A1 Overview
To pretrain E2IM v, we designed a Transformer-based autoencoder, consisting of voltage encoder $\mathcal{E}_{vol}$ and tactile decoder $\mathcal{D}_{tact}$. As shown in Fig. 2, the voltage autoencoder trains the input E2IM v in a self-supervised loop $\textbf{v}\rightarrow\hat{\textbf{v}}$. It first maps the input v to the latent representation $\textbf{z}_{v}$:
@@
$\arg\min_{\mathcal{E}_{vol},\mathcal{D}_{vol}}\mathcal{L}_{vol}=\|\hat{\textbf{v}}-\textbf{v}\|^{2}.$ (5)
This objective function aims to minimize the reconstruction error by optimizing the voltage encoder $\mathcal{E}_{vol}$ and decoder $\mathcal{D}_{vol}$.
@@ III-A2 Masked autoencoder framework
The pretraining model for E2IM utilizes the masked autoencoder framework [37] based on vision Transformer (ViT) [39]. The key element for this model is the Transformer layer [40], which processes data in sequential tokens. As described in Section II, voltage data v is represented by E2IM, a $64\times 64$ matrix. To process this form, E2IM data is pre-processed through masking and tokenization, as illustrated in Tokenization in Fig. 2. Specifically, for a batch of input E2IM matrices, $\textbf{v}\in\mathbb{R}^{64\times 64}$ is reshaped into a sequence of flattened 2D patches $\textbf{v}_{p}\in\mathbb{R}^{N\times(P^{2})}$. Here, $(P,P)=4^{2}$ is the resolution of each patch, $N=\frac{64\times 64}{4^{2}}=16\times 16$ is the resulting number of the patches.
Then, these patches $\textbf{v}_{p}$, preceded by an instructive class token $\textbf{v}_{cls}$ at the start of the sequence as $\textbf{v}_{p}^{0}$, are linearly projected into the Transformer layer and combined with patch and positional embeddings [39] for tokenization. The resulting input sequence, denoted as voltage token $\textbf{z}_{0}$, is computed as follows:
@@
The mask sampling strategy is applied to train the voltage encoder $\mathcal{E}_{vol}$. Specifically, a high masking ratio ($\alpha=75\%$ [37]) randomly masks $4\times 4$ $\textbf{v}_{p}\textbf{E}$ tokens globally. This approach maximizes information entropy and sets the masked tokens as prediction targets, enhancing the model’s learning generalizability. It leaves only the unmasked voltage tokens $(1-\alpha)\textbf{z}_{0}$ as input to the Transformer encoder $\mathcal{E}_{vol}$ to extract features:
@@
where $l=2\cdots(L+1)$ denotes the $l_{th}$ layer, $L=(1-\alpha)N$ represents the total number of unmasked voltage tokens $\textbf{v}_{p}$. Then the outputs are processed by multi-layer perception (MLP) as follows:
@@
To complete Equation 4, the voltage decoder $\mathcal{D}_{vol}(\textbf{z}_{v})$ utilizes positional embedding information $\textbf{E}_{pos}$ retained during tokenization to predict masked tokens and re-position them in the E2IM image, as shown in Fig. 2. The predicted masked region combines the unmasked part with the reconstructed E2IM $\hat{\textbf{v}}$. Unlike previous methods that processed voltage data v sequentially [27] and ignored the spatial arrangement of sensors, our approach employs a 2D E2IM matrix representation. This 2D form enables the PTET model to capture and save spatial information associated with electrodes in $\textbf{E}_{pos}$, ensuring the positional information is maintained throughout the learning process.

### v2

@@ Abstract
Tactile sensing is fundamental to robotic systems, enabling interactions through physical contact in multiple tasks. Despite its importance, achieving high-resolution, large-area tactile sensing remains challenging. Electrical Impedance Tomography (EIT) has emerged as a promising approach for large-area, distributed tactile sensing with minimal electrode requirements which can lend itself to addressing complex contact problems in robotics. However, existing EIT-based tactile reconstruction methods often suffer from high computational costs or depend on extensive annotated simulation datasets, hindering its viability in real-world settings. To address this shortcoming, here we propose a Pre-trained Transformer for EIT-based Tactile Reconstruction (PTET), a learning-based framework that bridges the simulation-to-reality gap by leveraging self-supervised pretraining on simulation data and fine-tuning with limited real-world data. In simulations, PTET requires 99.44% fewer annotated samples than equivalent state-of-the-art approaches (2,500 vs. 450,000 samples) while achieving reconstruction performance improvements of up to 43.57% under identical data conditions. Fine-tuning with real-world data further enables PTET to overcome discrepancies between simulated and experimental datasets, achieving superior reconstruction and detail recovery in practical scenarios. PTET’s improved reconstruction accuracy, data efficiency, and robustness in real-world tasks establish it as a scalable and practical solution for tactile sensing systems in robotics, especially for object handling and adaptive grasping under varying pressure conditions. All codes and data are publicly available in Edinburgh DataShare with the identifier https://doi.org/10.7488/ds/7985.
@@ I Introduction
Tactile sensing is a foundational technology in robotics, allowing robots to interact with their environment by perceiving pressure, texture, and force through physical contact[1, 2]. This capability is essential for tasks such as object manipulation, human-robot interaction, and environment exploration [3, 4]. Despite its significance, achieving high-resolution, large-area tactile sensing remains a challenge due to inherent trade-offs among accuracy, scalability, and complexity in sensor design. Current tactile sensors typically rely on capacitive[5, 6], piezoresistive[7], or piezoelectric[8] mechanisms. However, the spatial resolution of these systems is constrained by the number of sensor array pixels, which requires densely packed sensor arrays for higher spatial resolution [9, 10, 11]. This not only complicates fabrication and increases costs but also limits flexibility and scalability for expanded tactile sensing tasks [12].
@@
To overcome these limitations, Electrical Impedance Tomography (EIT) has emerged as a promising alternative [13]. EIT offers several advantages over traditional methods, including the ability to achieve large-area tactile sensing with fewer boundary electrodes[14, 15, 16]. By injecting small electrical currents through electrode pairs and measuring the resulting voltage changes, EIT reconstructs the conductivity distribution of a conductive soft layer to infer tactile inputs such as pressure or deformation [17]. However, EIT tactile reconstruction poses significant challenges in terms of accuracy and computational efficiency [18].
@@
Figure 1: Electrical Impedance Map (EIM). (a) Sensing principle. (b) Original measurements. (c) Construction of EIM: an example. (d) Enhanced EIM (E2IM).
Model-based methods[19, 20] have been typically employed for EIT reconstruction, solving the inverse problem non-iteratively or iteratively to estimate the conductivity distribution from voltage measurements. While these approaches do not require pre-collected datasets, they often produce suboptimal reconstruction results, especially for dynamic tactile sensing tasks. This limitation is largely attributed to EIT’s inherently ill-posed and ill-conditioned inverse problem [18], exacerbating reconstruction quality issues. Furthermore, the accuracy of model-based approaches is inherently limited due to simplified assumptions on the governing physical model, sensor geometry and materials properties [21].
Data-driven approaches, especially deep learning, have gained traction in EIT for mapping voltage measurements directly to tactile images, bypassing traditional iterative and linear methods [22]. While effective, these methods rely on large annotated datasets [23, 24, 25], which are impractical to collect at scale due to the labor-intensive nature of tactile data acquisition. As a result, training often depends on simulated data, which introduces domain gaps and reconstruction errors in real-world deployment, such as pressure inaccuracies, residual artefacts, and shape distortions [26].
To mitigate this, self-supervised pretraining has emerged as a powerful strategy, allowing models to learn generalizable features from unlabelled data and adapt to downstream tasks via fine-tuning on limited labelled samples. This paradigm has shown success in control [27], navigation [28], visual perception [29], and sim-to-real transfer [30]. Recent breakthroughs in masked prediction models, such as BERT [31] and MAE [32], further highlight its scalability for high-dimensional data.
Inspired by this, we propose a self-supervised EIT pretraining framework that learns from unlabelled real-world voltages and fine-tunes on limited real voltage–tactile pairs. This enables a real-to-real learning pipeline that avoids heavy reliance on simulation and reduces domain shift errors in tactile reconstruction.
@@
In this paper, we propose PTET (Pre-trained Transformer for EIT-based Tactile Reconstruction), a novel framework that tackles two core challenges in learning-based EIT tactile sensing: (1) reliance on large annotated datasets and (2) simulation-to-reality gaps. PTET employs self-supervised pretraining on unlabelled voltage data and few-shot fine-tuning on limited real-world pairs, enabling accurate tactile reconstruction with minimal annotation and reduced simulation-induced errors. This dual-phase learning strategy allows PTET to bridge the gap between experimental conditions and practical applications. Unlike traditional learning-based approaches that require extensive labelled datasets from simulation, PTET effectively operates with a small fraction of the real-world data, making it scalable and practical for real-world deployments. Our main contributions are:
@@
  • We develop a large-area, flexible EIT tactile sensor using a dual-conductivity layer: hydrogel for biocompatibility and a carbon black–graphite (CBG) composite for enhanced conductivity and flexibility. The sensor achieves high sensitivity and detects compressive strains as small as 1 mm.
  • We demonstrate the first high-resolution, learning-based tactile reconstruction trained solely on limited experimental data. This real-to-real learning pipeline eliminates simulation-related errors, advancing the practicality of EIT tactile systems.
@@ II Enhanced Electrical Impedance Map
In each EIT measurement, current is injected through a pair of adjacent electrodes, and voltages are measured between neighboring pairs among the remaining electrodes, excluding the injectors (see Fig. 1a). Specifically, for a 16-electrode setup, current is injected between 16 adjacent electrode pairs, and for each excitation, voltages are measured across the remaining 13 pairs (excluding the injectors), yielding 208 measurements (16 $\times$ 13). Due to the reciprocity theorem and adjacent measurement protocol, only 104 of these are independent [33].
Figure 2: Overview of PTET. Our method comprises three main processes: bottom— 1) Pretraining with E2IM, for an effective and robust voltage encoder $\mathcal{E}_{vol}$, through self-supervised learning by predicting masked regions in grey. 2) Pretraining with tactile map, through self-supervised learning, for a robust tactile decoder $\mathcal{D}_{tact}$; top— 3) Fine-tuning with limited pairs from real-world measurements, the fine-tuned model will be deployed for reconstruction (*: optional in fine-tuning). Notably, during fine-tuning, $\mathcal{E}_{vol}$’s parameters are frozen, while $\mathcal{D}_{tact}$’s are tunable.
[…]
@@ III PTET: Pre-trained Transformer for EIT-based Tactile Reconstruction
The reconstruction of tactile representations t from EIT voltage measurements v, represented in the format of E2IM, can be described by a direct mapping: $f:\textbf{v}\rightarrow\textbf{t}$.
@@
    Pretraining with E2IM: In this stage, a masked autoencoder (MAE) $\mathcal{D}_{vol}(\mathcal{E}_{vol}(\textbf{v}))$ is employed to extract features from unlabeled voltage data v, which is represented by E2IM. This process trains an effective and robust encoder $\mathcal{E}_{vol}(\textbf{v})$, functioning mapping $p_{v}:\textbf{v}\rightarrow\textbf{z}_{v}$, which compresses the input v to a latent variable $\textbf{z}_{v}$;
@@
PTET adopts a dual-phase training scheme: a pretraining phase with 1) and 2), and a fine-tuning phase with 3). In the pretraining phase, 1) the voltage encoder $\mathcal{E}_{\text{vol}}$ is trained on large-scale unannotated E2IM data, and 2) the tactile decoder $\mathcal{D}_{\text{tact}}$ is trained separately on unannotated tactile maps. This allows both components to learn generalizable features from readily available simulation data. In the fine-tuning phase, 3) $\mathcal{E}_{\text{vol}}$ and $\mathcal{D}_{\text{tact}}$ are jointly refined using limited paired real-world data to align the voltage and tactile domains. This enables accurate tactile reconstruction from real E2IM inputs. PTET employs an asymmetric encoder–decoder architecture in all stages to improve feature representation efficiency, as detailed in the following sections.
@@ III-A1 Overview
To pretrain E2IM v, we designed a Transformer-based autoencoder, consisting of tactile encoder $\mathcal{E}_{vol}$ and tactile decoder $\mathcal{D}_{tact}$. As shown in Fig. 2, the voltage autoencoder trains the input E2IM v in a self-supervised loop $\textbf{v}\rightarrow\hat{\textbf{v}}$. It first maps the input v to the latent representation $\textbf{z}_{v}$:
@@
$\arg\min_{\mathcal{E}_{vol},\mathcal{D}_{vol}}\mathcal{L}_{vol}=\mathbb{E}_{\textbf{v}}\left[\|\hat{\textbf{v}}-\textbf{v}\|^{2}\right].$ (4)
This objective function aims to minimize the expectations ($\mathbb{E}$) of reconstruction error by optimizing the voltage encoder $\mathcal{E}_{vol}$ and decoder $\mathcal{D}_{vol}$.
@@ III-A2 Masked autoencoder framework
The pretraining model for E2IM utilizes the masked autoencoder framework [32] based on vision Transformer [34]. The key element for this model is the Transformer layer [35], which processes data in sequential tokens. As described in Section II, voltage data v is represented by E2IM, a $64\times 64$ matrix. To process this form, E2IM data is pre-processed through masking and tokenization, as illustrated in Tokenization in Fig. 2. Specifically, for a batch of input E2IM matrices, $\textbf{v}\in\mathbb{R}^{64\times 64}$ is reshaped into a sequence of flattened 2D patches $\textbf{v}_{p}\in\mathbb{R}^{N\times(P^{2})}$. Here, $(P,P)=4^{2}$ is the resolution of each patch, $N=\frac{64\times 64}{4^{2}}=16\times 16$ is the resulting number of the patches.
Then, these patches $\textbf{v}_{p}$, preceded by an instructive class token $\textbf{v}_{cls}$ at the start of the sequence as $\textbf{v}_{p}^{0}$, are linearly projected into the Transformer layer and combined with patch and positional embeddings [34]. The resultant input voltage token $\textbf{z}_{0}$, is computed as follows:
@@
We apply a mask sampling strategy to train the voltage encoder $\mathcal{E}_{\text{vol}}$. Following [32], a high masking ratio ($\alpha=75\%$) randomly masks $4\times 4$ $\textbf{v}_{p}\textbf{E}$ tokens globally, maximizing information entropy and setting masked tokens as prediction targets. Only the remaining unmasked tokens $(1-\alpha)\textbf{z}_{0}$ are fed into the Transformer encoder $\mathcal{E}_{\text{vol}}$ for feature extraction:
@@
where $l=2\cdots(L+1)$ denotes the $l_{th}$ layer, $L=(1-\alpha)N$ represents the total number of unmasked voltage tokens $\textbf{v}_{p}$. Then the outputs are processed by multi-layer perception as follows:
@@
To complete Equation (3), the voltage decoder $\mathcal{D}_{\text{vol}}(\mathbf{z}_{v})$ uses positional embeddings $\mathbf{E}_{pos}$ retained from tokenization to predict masked tokens and reposition them within the E2IM image (Fig. 2). The final reconstruction $\hat{\mathbf{v}}$ combines these predictions with the unmasked regions. Unlike prior methods that process voltage data $\mathbf{v}$ sequentially [23] and ignore sensor spatial layout, our 2D E2IM representation preserves electrode spatial information via $\mathbf{E}_{pos}$, maintaining positional context throughout learning. As shown in Table I, increasing the dimension and size of $\mathbf{E}_{pos}$ with E2IM significantly improved the feature extraction capability of $\mathcal{E}_{vol}$.